APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Wednesday October 20, 2004 (MJD 53299) ---=== OBSERVING TEAM ===--- Swing: Pete Newman Night: Steph Snedden, Atusko Nitta Visitors: Fritz de Jongh (SNe Team, FNAL) All Mighty: Hercules ---=== OBSERVING PLAN ===--- Science. ---=== OBSERVING SUMMARY ===--- o We finished plate 1961, 1634, 1782. o We got signal on 1910 o We could not plot the spectra using apoplot for the data on 1634. We don't know why. o Red Sox won. o Dvorzak seems to help get rid of clouds. ---=== OBSERVING LOG ===--- Swing shift =========== The focus positions of the spectrographs were checked and adjusted. See end of log for details. Night ======= 00:30Z Opened. 01:00Z Cartridge 3 Plate 1961. M2 (tempted by coffee plate) guideStars on gotoField. Seeing is 1.4", but sky is not photometric. During the second exposure, the sky became photometric. The moon is bright, effecting our efficiency. DIMM required us to restart its server. All was well after that. SoS recommended postCalib and we obtained them. The sky became non photometric again.... 03:06Z cart 7 plate 1910 No guide stars with or without the offsets from the previous plate. Found fk5 when we went -10" in dec. The second exposure did not make the s/n^2 of 2. The clouds are thickening. Exp 28962 declared bad. 04:50Z Closed due to clouds. 08:45Z As Steph started to play music by Dvorzak, the sky started to clear. We mounted another cartridge inside the enclosure since the sky was clearing. Unfortunately, the wind started to pick up (which is probably why the clouds were moving away fast). So we wait... 09:15Z The wind is high, but not too high. So we open with cart 1 plate 1634. Guide stars on gotoField.1.3" seeing. With gust at 30m/h, DIMM will not be able to estimate the seeing reliably (the telescope shakes too much in strong wind0. So we did not open DIMM. In the excitement of opening and getting photons again, we started the first science exposure before the preCalib. We take calib between the first and second science exposures. 09:50Z The cloud camera says the sky is photometric. We hope the clouds we see in the IR camera looking at the telescoe pwill stay where they are. 10:20Z Although the std dev is still low, we see the clouds coming in from the south where we are pointing at. Stay away Clouds! Combined with previous exposures, this plate was declared done. We tried to plot some science spectra using apoplot for plate 1634. Although we were able to do apoplotsky for this plate, we could not apoplot to plot any science spectra. The command failed with the following message. IDL 5.5a> apoplot, 1634, 439, nsmooth=5 APOPLOT: Selecting MJD=53299 (override this with MJD keyword) MRDFITS: File access error % Expression must be a structure in this context: TSET. % Execution halted at: TRACESET2XY 47 /local/ups/prd/Linux/idlutils/v4_10_7/pro/trace/traceset2xy.pro % APOPLOT1 182 /local/ups/prd/Linux/idlspec2d/v4_10_7/pro/apo2d/apoplot.pro % APOPLOT 447 /local/ups/prd/Linux/idlspec2d/v4_10_7/pro/apo2d/apoplot.pro % TRACESET2XY 47 /local/ups/prd/Linux/idlutils/v4_10_7/pro/trace/traceset2xy.pro % APOPLOT1 182 /local/ups/prd/Linux/idlspec2d/v4_10_7/pro/apo2d/apoplot.pro % APOPLOT 447 /local/ups/prd/Linux/idlspec2d/v4_10_7/pro/apo2d/apoplot.pro % TRACESET2XY 47 /local/ups/prd/Linux/idlutils/v4_10_7/pro/trace/traceset2xy.pro % APOPLOT1 182 /local/ups/prd/Linux/idlspec2d/v4_10_7/pro/apo2d/apoplot.pro % APOPLOT 447 /local/ups/prd/Linux/idlspec2d/v4_10_7/pro/apo2d/apoplot.pro % $MAIN$ Just in case this was due to this plate having many test exposures, we decided to try running the command with exposure number specifide. Specifying exposure number did not help either. 10:35Z Next plate cart 9 plate 1782. We got fiducial too large to correct in Alt as we slewed to the field and the following fiducials all were "too large to correct". We saw no guide stars. We forced fiducial and then reslewed to the field and found guide stars. We can do apoplot on this plate to see the science spectra. Anything wrong with the previous plate exposures?? THe clouds skirt around us. Rarely ever going above 30 deg., according to the cloud camera. 12:00Z Closed. endNight finished! ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 4906 22:02Z 23:03Z 100 O 69.45 84.65 115 ignore - on telescope, but enclosure not darkened. ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n):y QA Procedures Done (y/n):y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 28939, plate -9999 ------- 22:03 28939 0.0 bias ----- sequence 28940, plate 1634 ------- 23:22 28940 0.0 bias 23:28 28941 2.1 arc 23:30 28942 2.1 arc 23:40 28943 2.1 arc 23:43 28944 2.1 arc 23:58 28945 2.1 arc 00:00 28946 2.1 arc ----- sequence 28947, plate -9999 ------- 00:08 28947 0.0 bias 00:11 28948 0.0 bias ----- sequence 28949, plate 1961 ------- 17.4 28.3 16.6 39.9 00:19 28949 10.0 flat ----- sequence 28949, plate 1961 ------- 3.5 7.5 6.2 9.2 00:22 28950 2.0 arc 01:23 28951 10.0 flat 01:25 28952 2.0 arc 01:46 28953 900.1 target 02:10 28954 1200.1 target 02:32 28955 1200.1 target 02:50 28956 900.1 target 02:55 28957 10.0 flat 02:57 28958 2.0 arc ----- sequence 28959, plate 1910 ------- 3.5 7.5 6.2 9.2 03:15 28959 10.0 flat 03:18 28960 2.0 arc 03:37 28961 900.1 target 04:02 28962 694.9 target ----- sequence 28963, plate 1634 ------- 17.8 10.7 20.3 12.3 (done combined 09:52 28963 1200.1 target with previous exposures) 09:56 28964 10.0 flat 09:58 28965 2.0 arc 10:24 28966 1200.1 target ----- sequence 28967, plate 1782 ------- 34.6 20.2 30.9 19.9 10:41 28967 10.0 flat 10:43 28968 2.0 arc 11:07 28969 1200.1 target 11:30 28970 1200.1 target 11:48 28971 900.1 target 11:51 28972 10.0 flat 11:54 28973 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 01:29Z 3 1961 29.45 0.0021 53.30 -0.0021 204.31 0.0088 1.000200 03:12Z 7 1910 18.09 0.0006 46.17 -0.0043 195.22 -0.0114 1.000120 09:31Z 1 1634 -14.95 0.0018 50.14 -0.0031 167.31 -0.0179 1.000440 10:44Z 9 1782 139.96 0.0012 54.53 -0.0030 -68.81 -0.0068 1.000330 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL7975 Stays: JL7982 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 01:29Z 3 1961 1.00020 -1974 -389 75 29 53.3 9.4 18 224 gdr 1.4 03:12Z 7 1910 1.00012 -1184 163 -50 18 46.2 10.3 20 228 gdr 1.3 09:31Z 1 1634 1.00044 -4344 -2441 -25 -15 50.1 9.1 25 227 gdr 1.3 10:46Z 9 1782 1.00033 -3258 -1555 -25 140 54.7 8.7 18 227 gdr 1.2 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 21:37Z 57 22 4 282 (WNW) 372 - 22:10Z 54 24 22 250 (WSW) 518 - 22:46Z 54 25 20 227 (SW) 530 - 23:18Z 53 23 19 239 (WSW) 439 - 23:50Z 51 24 12 220 (SW) 506 - 00:22Z 49 25 15 218 (SW) 581 - 01:03Z 48 27 11 224 (SW) 629 - 01:34Z 49 27 10 234 (SW) 661 - 02:07Z 49 27 11 223 (SW) 689 - 02:40Z 50 27 13 239 (WSW) 758 - 03:12Z 50 26 19 228 (SW) 802 - 03:43Z 49 27 20 211 (SSW) 786 - 04:15Z 49 28 22 207 (SSW) 809 - 04:47Z 50 27 23 203 (SSW) 778 - 05:20Z 50 27 27 218 (SW) 842 - 05:52Z 50 27 23 220 (SW) 770 - 06:22Z 49 27 23 213 (SSW) 865 - 06:52Z 49 26 21 199 (SSW) 677 - 07:25Z 49 27 21 202 (SSW) 895 - 07:57Z 48 27 26 216 (SW) 785 - 08:29Z 48 28 20 212 (SSW) 634 - 08:59Z 48 27 28 224 (SW) 581 - 09:31Z 48 27 25 226 (SW) 601 - 10:03Z 48 28 22 219 (SW) 560 - 10:33Z 47 27 15 213 (SSW) 468 - 11:04Z 48 28 15 235 (SW) 453 - 11:35Z 47 26 12 262 (W) 467 - 12:05Z 46 22 15 261 (W) 343 - ---=== TELESCOPE STATUS ===--- 00:30Z Opened! 04:50Z Enclosure on, but fans on and doors open. 08:45Z 12:00Z Closed and fans off. Telescope at stow with spectro auto-fill connected. No cartridge on telescope. Only the corrector lens. No interlock bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_142_0 Watcher: v2_29_0 MCP: v5_25_0 TPM: tpm_v2_42_0 AstroDa: v14_47 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v1_94 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5220 0.5250 Axial B 0.6020 0.6020 Axial C 0.9070 0.9050 Trans D -8.6200 -8.6120 Lateral E 10.0330 10.0203 Lateral F 11.5951 11.5951 GALILS Commanded: 4800. -2150. 2600. -6900. -7850. -7900. Actual: 4782. -2150. 2618. -6880. -7862. -7902. SETMIR VALUES PriDesOrient: 0.00 -2.76 1.04 57.00 151.01 PriOrient: 0.00 -3.04 0.88 57.42 149.40 SECONDARY: ---------- Focus: 0.00 Air Temp.: 13.1 Alt.: 29.998954 MIGS TONIGHT NOMINAL Axial A 1.5410 1.5390 Axial B 1.0480 1.0440 Axial C 1.1590 1.1590 Trans D -0.2140 -0.2020 GALILS Commanded: 1619145. 1604991. 1528443. -13800. -9950. Actual: 1619188. 1604973. 1528571. -13820. -9964. SETMIR VALUES SecDesOrient: 1257.00 27.97 -10.00 -90.00 257.89 SecOrient: 1256.96 28.01 -9.99 -89.93 257.47 ---=== PROBLEMS IN DETAIL ===--- ---=== SPECTROGRAPHS' FOCUS POSITIONS ===--- Initial positions: sp1: Coll_motor_A -8060 Coll_motor_B -7097 Coll_motor_C -8760 sp2: Coll_motor_A -11804 Coll_motor_B -6481 Coll_motor_C -6702 sp1> loadCart 1 1634 sp1> date Wed Oct 20 23:25:56 UTC 2004 sp1> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 9.49786 tpm_TM_TelairAvg.VALA 11.105 TCC> track 121,30 mount /rotan=45 /rotty=mount sp1> goSpecFocus IDL> collimate, 28941 COLLIMATE: Mean offset = 0.09 pix COLLIMATE: Camera b1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of -3.3 degrees for spectro-1 COLLIMATE: Mean offset = -0.24 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of 8.5 degrees for spectro-2 COLLIMATE: Mean offset = -0.20 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 1815 steps for spectro-1 COLLIMATE: Mean offset = 0.11 pix COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of -964 steps for spectro-2 sp1> mechPiston 1815 sp2> mechPiston -964 sp1> goSpecFocus IDL> collimate, 28943 COLLIMATE: Mean offset = -0.04 pix COLLIMATE: Camera b1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 1.3 degrees for spectro-1 COLLIMATE: Mean offset = -0.14 pix COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 5.0 degrees for spectro-2 COLLIMATE: Mean offset = -0.04 pix COLLIMATE: Camera r1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 398 steps for spectro-1 COLLIMATE: Mean offset = 0.04 pix COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of -344 steps for spectro-2 Since b2 is so close to the edge of tolerance and the other cameras well centered, I adjusted the b2 camera focus ring by +5 deg. Unfortunately, in releasing the focus locking ring, the focus ring also moved by an unknown amount in the counterclockwise (out) direction, although I was fairly certain where it had started (+3 deg), so I moved it to +8 deg (via a large negative move to remove backlash. sp1> goSpecFocus IDL> collimate, 28945, docams="b2" COLLIMATE: Mean offset = -0.00 pix COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 0.2 degrees for spectro-2 I got it right! Final mirror positions: sp1: Coll_motor_A -6244 Coll_motor_B -5271 Coll_motor_C -6935 sp2: Coll_motor_A -12769 Coll_motor_B -7444 Coll_motor_C -7666 Too Large to correct fiducials ========================== While going to field on cart3: alt fiducial 59.12 deg correction for altitude -640 is too large But the following crossing were fine. While going to cart 7 field, altitude (correction for altitude -640 is too large (2004-10-21 02:57:19Z sdssmcp): 2004-10-21 02:57:29Z ) but the fiducial crossing following the above was fine. While doing a instCHange inside the enclosure, we got the following. altitude (correction for altitude -672 is too large (2004-10-21 08:55:50Z sdssmcp): 2004-10-21 08:55:58Z ) The following altitude corssing was fine. So no ill effect. When going to field on cart 9, we got the too large to correct message and we did have to accept the fiducial since all the fiducial following this one was too large to correct. altitude (correction for altitude 44400 is too large (2004-10-21 10:31:36Z sdssmcp): 2004-10-21 10:31:47Z ) Max Acceleration errors ========================== When we were closing due to clouds, we got: altitude (calc_frames: Max accl. for altitude exceeded: -125/1000 (2004-10-21 04:54:17Z sdssmcp): 2004-10-21 04:54:25Z ) TPM errors =========== dewar (Imager Dewar Scale Communications Fault:: 2004-10-21 00:09:54Z) m2 mig (Communications Fault: 2004-10-21 00:08:19Z)